### v1

@@ 4 Theory
for all $1\leq i\leq r$, $1\leq k\leq K$, where $C_{0,K}$ is a constant depending on $K$ only and $C$ is a positive numeric constant.
Again, we assume $T\leq C\exp\left(d_{\min}^{\vartheta/(2\vartheta+4)}\right)$ and $T\leq C\exp\left((d_{\min}/r)^{\gamma_{1}/2}\right)$ primarily for simplifying the convergence rate. It is important to note that the error bound $\psi_{0}$ for initialization is intended for each individual loading vector $a_{ik}$. When applying Algorithm 3, which requires the inverse of $\widehat{A}_{k}^{\top}\widehat{A}_{k}$, the condition $\sqrt{r}\psi_{0}\lesssim 1$ ensures a reliable initial estimate of the loading matrix $\widehat{A}_{k}$. Although the other components in (23) may seem complex, they are designed to ensure the error contraction effect in each iteration. This ensures that as iterations progress, the error bound will approach the desired statistical upper bound.
@@
$\|\widehat{a}_{ik}^{{\rm iso}}\widehat{a}_{ik}^{{\rm iso}\top}-a_{ik}a_{ik}^{\top}\|_{2}=O_{\mathbb{P}}\left(\sqrt{\frac{d_{\max}}{\lambda_{r}T}}+\frac{1}{\lambda_{r}}\right).$ (25)
In the typical strong factor models where $\lambda_{1}\asymp\lambda_{r}\asymp d$, the rate simplifies to $O_{\mathbb{P}}(\sqrt{d_{\max}/(dT)}+1/d)$. This rate is significantly faster than that found in the vector factor model.
@@
$\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(a_{ik}^{\top}\widehat{a}_{ik}^{{\rm iso}})a_{ik}=P_{a_{ik},\perp}\left[\frac{1}{\Theta_{ii}T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash{k}}b_{i\ell}^{\top}\right)\right]+O_{\mathbb{P}}\left(\frac{1}{\Theta_{ii}}\left(\frac{d_{k}}{T}+\sqrt{\frac{d_{k}}{T}}+1\right)\right),$
where $P_{a_{ik},\perp}=I_{d_{k}}-a_{ik}a_{ik}^{\top}$ and $\Theta=(\Theta_{ij})_{r\times r}$, where $\Theta$ is defined in Assumption 4.2. This enables the determination of asymptotic distributions for linear forms of $a_{ik}$.
The following theorem shows the asymptotic distribution of a linear form of the factor loading vector $u^{\top}a_{ik}$ for some fixed vector $u$. Note that in the strong factor model, we have $\Theta_{ii}\asymp d$ for all $1\leq i\leq r$.
@@
Suppose the conditions in Theorem 4.2 are satisfied. Let $\lambda_{1}\asymp\lambda_{r}$. Assume that $\lim\inf_{d_{k}\to\infty}$ $\|P_{a_{ik},\perp}u\|_{2}>0$, for each $1\leq i\leq r,1\leq k\leq K$, we have:
@@
    If $T/(d_{k}\Theta_{ii})\to 0$, then
@@
    where $\sigma_{u,ik}^{2}=h_{ik}^{\top}\Sigma_{e}h_{ik}/\Theta_{ii}$, $h_{ik}=b_{iK}\odot\cdots\odot b_{i,k+1}\odot P_{a_{ik},\perp}u\odot b_{i,k-1}\odot\cdots\odot b_{i1}\in\mathbb{R}^{d}$ and $\odot$ represents Kronecker product.
  2. (ii)
    If $d_{k}\Theta_{ii}=O(T)$, then
In Theorem 4.3, we focus on vectors $u$ with the property that $\|P_{a_{ik},\perp}u\|_{2}>0$ when $d_{k}$ is large, effectively presupposing that $\sin\angle(u,a_{ik})>0$. Conversely, when $\sin\angle(u,a_{ik})=0$, the convergence rate of the estimated linear form is faster, and its asymptotic distribution is a blend of $\chi_{1}^{2}$ distributions.
@@
Suppose the conditions in Theorem 4.2 are satisfied. Let $\lambda_{1}\asymp\lambda_{r}$. For each $1\leq i\leq r,1\leq k\leq K$, we have:
@@
    If $T/(d_{k}\Theta_{ii})\to 0$, then
[…]
    where $\varpi_{j},1\leq j\leq d_{k}$ are the eigenvalues of $\Sigma_{ik}^{1/2}P_{a_{ik},\perp}\Sigma_{ik}^{1/2}$, with $\Sigma_{ik}=\Theta_{ii}^{-1}\mathbb{E}[({\cal E}_{t}\times_{\ell\neq k}^{K}b_{i\ell})({\cal E}_{t}\times_{\ell\neq k}^{K}b_{i\ell})^{\top}]$.
@@ 5 Simulation
In this section, we conduct empirical comparisons among different methods for estimating loading vectors across various simulation scenarios, and verify the limiting distribution of the estimated loading vectors. We assess the performance of the contemporary covariance-based iterative simultaneous orthogonalization procedure (CC-ISO) proposed in this paper, auto-covariance-based iterative simultaneous orthogonalization procedure by Han et al., (2023) (AC-ISO), and tensor principle component analysis (TPCA) by Babii et al., (2023). The auto-covariance considered by Han et al., (2023) is defined by the following lagged-cross product operator:
@@
$\max_{i,k}\|\widehat{a}_{ik}\widehat{a}_{ik}^{\top}-a_{ik}a_{ik}^{\top}\|_{2}.$
Throughout our analysis, the observations ${\cal Y}_{t}$’s are simulated according to model (1) with $K=2$. The true loading vectors are generated as follows: The elements of matrices $\widetilde{A}_{k}=\left(\widetilde{a}_{1k},\ldots,\widetilde{a}_{rk}\right)\in$ $\mathbb{R}^{d_{k}\times r},1\leqslant k\leqslant K$, are drawn from i.i.d. $N(0,1)$ distributions and then orthonormalized via QR decomposition. If $\delta=0$, set $A_{k}=\widetilde{A}_{k}$; otherwise, set $a_{1k}=\widetilde{a}_{1k}$ and $a_{ik}=\left(\widetilde{a}_{1k}+\theta\widetilde{a}_{ik}\right)/\left\|\widetilde{a}_{1k}+\theta\widetilde{a}_{ik}\right\|_{2}$ for all $i\geqslant 2$ and $1\leqslant k\leqslant K$, with $\vartheta=\delta/(r-1)$ and $\theta=\left(\vartheta^{-2/K}-1\right)^{1/2}$. In our simulation study, we vary the correlations between loading vectors through $\delta$. It is evident that an increase in $\delta$ leads to a higher degree of linear dependence among the loading vectors.
The factor processes $f_{it}$ exhibit weak temporal dependence and are generated as an independent AR(1) process multiplied by a scalar depending on $d_{1},d_{2}$ and $r$:
$f_{it}=w_{i}g_{it},$ (31)
where
@@
  • $w_{i}=\frac{1}{5}\times(r-i+1)\sqrt{d_{1}\times d_{2}}$.
In this case, the tensor factor model in (1) represents a typical strong factor model, where $\lambda_{i}=w_{i}^{2}=(r-i+1)^{2}d_{1}d_{2}/25=O(d)$ when $r$ is fixed. In Appendix C, we provide the result for $\phi=0.5$.
@@
    (Orthogonal loading matrix) Set $\delta=0$ so that the columns of loading matrix $A_{k}$ are orthonormal. Each entry of error term ${\cal E}_{t}$ is generated independently from $N(0,1)$.
@@
    (Non-orthogonal loading matrix) Vary $\delta$ in the set $\{0.1,0.3,0.5\}$ so that the columns of loading matrix $A_{k}$ are not orthogonal. Each entry of error term ${\cal E}_{t}$ is generated independently from $N(0,1)$.
@@
    (Serial correlation in ${\cal E}_{t}$) Set $\delta=0.2$. The errors ${\cal E}_{t}$ are generated according to ${\cal E}_{t}=\Psi_{1}^{1/2}Z_{t}\Psi_{2}^{1/2}$, where
@@
    (Weak factors) Set $r=3$, and $\delta=0.2$. The error terms are generated according to ${\cal E}_{t}=\Psi_{1}^{1/2}Z_{t}\Psi_{2}^{1/2}$, where
@@
    The scaling multiplier in factor process $w_{i}=(r-i+1)\times(d_{1}d_{2})^{1/\alpha}$, where $\alpha$ varies in the set $\{2.5,3,3.5,4\}$. Note that when $\alpha=2$, the factor structure is considered strong. A larger $\alpha$ indicates a weaker factor structure.
@@
In Configuration II, we assess the impact of non-orthogonal factor loadings on estimations using ISO algorithms and TPCA algorithm. Figure 3 shows the ratio of the estimation errors of CC-ISO to AC-ISO (first panel) and to TPCA (second panel) across different values of $\delta$ and dimensions. The error ratio of CC-ISO to AC-ISO remains around 0.15, indicating the superior accuracy of CC-ISO. The ratio remains relatively stable because the signal part in AC-ISO, albeit small, also increases with dimensions, resulting in limited improvements on the estimation. However, in the second panel, the error ratio of CC-ISO to TPCA converges as dimensions increase. This is because TPCA cannot identify non-orthogonal factor loading vectors, leading to stable estimation errors across varying dimensions. In contrast, CC-ISO successfully identifies non-orthogonal loading vectors, resulting in estimation errors converging to 0.
@@
Figure 4 shows the ratio of estimation errors of CC-ISO to AC-ISO under configuration III, designed to evaluate the robustness of proposed CC-ISO algorithm against serial correlation in the error term. We observe that CC-ISO’s performance improves monotonically as $T$ increases. In contrast, AC-ISO’s performance deteriorates as the serial correlations in the error term strengthens. This decline is due to the contamination of signals in the auto-covariance by the serial correlations in the error terms. However, CC-ISO demonstrates robustness against such serial correlations.
@@
We also examine the performance of Randomized Projection (RP) from Procedure 2 and compare it with Composite PCA (C-PCA), which corresponds to Algorithm 1 but without the steps for detecting close eigenvalues (Step 3, 5, 6, and 7 in Algorithm 1).
@@
    (C-PCA vs. RP-PCA) $r=5$. $d_{1}=d_{2}=\bar{d}$ with $\bar{d}\in\{20,40,80\}$ and $T\in\{100,200,500\}$. The columns of factor loadings $A_{k}$ are orthonormal and are generated as described in Configuration I. Furthermore, the factors $f_{it}$ are also orthonormal, generated using QR decomposition after deriving from AR(1) processes. In this setting, the singular values of the common components $\sum_{i=1}^{r}w_{i}g_{it}a_{i1}\otimes a_{i2}$ are solely determined by $w_{i}$. We set $w_{i}=w=10$ to ensure identical eigenvalues of common components. Error terms are generated from i.i.d. $N(0,1)$. Though the top $r$ eigenvalues of the $\widetilde{\Sigma}$ are not identical due to noise, their differences are relatively small, allowing randomized projection algorithms to ensure the accuracy of initial estimations. For the remaining parameters, we set $\nu=0.8$, $c_{0}=0.1$ and $L=2r^{2}$.
@@
Given the close empirical performances of CC-ISO under both initialization methods under configuration V, our focus shifts to the estimation errors of the initial estimations, as illustrated in Figure 6. RP algorithm outperforms the RC-PCA algorithm in terms of the accuracy of initial estimations, particularly pronounced when $d$ is smaller and $T$ is larger. This occurs because the sample covariance of $\operatorname{Vec}({\cal E}_{t})$ approaches the identity matrix as $d$ decreases and $T$ increases. Consequently, $\widetilde{\Sigma}$ is more likely to have eigenvalues that are closer together.
@@
where ${\cal F}_{t}\in\mathbb{R}^{r\times r}$ is the factor process in the Tucker factor model. In the CP factor model, ${\cal F}_{t}$ is diagonal with the $i^{th}$ diagonal element equal to $f_{it}$. In the mis-specification setting, we allow the off-diagonal entries to deviate from 0. Denote the $(i,j)^{th}$ entry of ${\cal F}_{t}$ by $f_{ijt}$. Let $f_{ijt}=w_{ijt}g_{ijt}$, where $g_{ijt}$ is generated as specified in (31). The configuration is as follows:
@@
    (Mis-specification) $r=3,(d_{1},d_{2})\in\{(40,40),(40,60),(60,60)\}$ and $T\in\{100,300,500\}$. The loading vectors and error terms are generated as in Configuration IV, allowing for correlation between loading vectors and weak cross-sectional correlation in the error term. $w_{ijt}=\sqrt{d_{1}d_{2}}/5$ if $i=j$ and $w_{ijt}=(d_{1}d_{2})^{1/\alpha}/5$ with $\alpha\in\{3,4,5\}$. A smaller $\alpha$ indicates a more severe mis-specification in the model.

### v2

@@ 4 Theory
for all $1\leq i\leq r$, $1\leq k\leq K$, where $C_{0,K}$ and $C_{1,K}$ are some constants depending on $K$ only and $C$ is a positive numeric constant.
It is important to note that the error bound $\psi_{0}$ for initialization is intended for each individual loading vector $a_{ik}$. When applying Algorithm 3, which requires the inverse of $\widehat{A}_{k}^{\top}\widehat{A}_{k}$, the condition $\sqrt{r}\psi_{0}\lesssim 1$ ensures a reliable initial estimate of the loading matrix $\widehat{A}_{k}$. Although the other components in (23) may seem complex, they are designed to ensure the error contraction effect in each iteration. This ensures that as iterations progress, the error bound will approach the desired statistical upper bound. A more detailed discussion of initial estimates can be found in Appendix E.
@@
$\max_{i}\|\widehat{a}_{ik}^{{\rm iso}}\widehat{a}_{ik}^{{\rm iso}\top}-a_{ik}a_{ik}^{\top}\|_{2}=O_{\mathbb{P}}\left(\sqrt{\frac{d_{k}}{\lambda_{r}T}}+\frac{1}{\lambda_{r}}\right).$ (25)
In the typical strong factor models where $\lambda_{1}\asymp\lambda_{r}\asymp d$ (i.e. $w_{i}\asymp\sqrt{d}$), the rate simplifies to $O_{\mathbb{P}}(\sqrt{d_{k}/(dT)}+1/d)$. This rate is significantly faster than that found in the vector factor model. Moreover, in comparison with the initial estimator discussed in Remark 4.1, the term $1/d$ arises from the cross-sectional dependence among the noise tensor ${\cal E}_{t}$ in (1), and is therefore irreducible. Notably, the bias $\delta$ is eliminated by the iterative refinement process. More importantly, the main source of estimation error is reduced from $O_{\mathbb{P}}(\sqrt{1/T})$ to $O_{\mathbb{P}}(\sqrt{d_{k}/(dT)})$.
Remark 4.4.
To further examine the constant in the error contraction condition (23), we assume that the constant in the convergence rate of the final estimator (24) satisfies $C_{0,K}=C_{0}\alpha^{2-2K}$, for some absolute constant $C_{0}>0$, and define $\alpha=\sqrt{1-\delta_{\max}}-(r^{1/2}+1)\psi_{0}/\sqrt{1-1/(4r)}$. Then, condition (23) can be detailed as follows:
$\alpha>0,\ C_{0}\alpha^{2-2K}\left(\frac{\lambda_{1}}{\lambda_{r}}\right)\psi_{0}^{2K-3}<1,$
$C_{0}\alpha^{2-2K}\sqrt{\frac{\lambda_{1}}{\lambda_{r}}}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)\psi_{0}^{K-2}<1.$
Assuming $\lambda_{1}\asymp\lambda_{2}\cdots\asymp\lambda_{r}$, condition (23) requires that $r\psi_{0}\lesssim 1$, which in turn implies $r\delta\lesssim 1$ under the setting of Theorem 4.1(i). This condition is automatically satisfied under orthogonal loadings, but it also accommodates the case with mild non-orthogonal loadings.
The following Theorem 4.3 specifies the convergence rate for the estimated factors $f_{it}$ [^10] Asymptotic normality results for estimated factors are developed in our companion paper on diffusion index forecasting: https://papers.ssrn.com/sol3/papers.cfm?abstract_id=5213594, which builds directly on the theoretical framework established here..
Theorem 4.3.
Suppose Assumptions 4.1, 4.2, 4.3 hold. Assume that $\delta_{\max}=\max_{k\leq K}\delta_{k}<1$ with $\delta_{k}$ defined in (16), $r=O(T)$, $1\lesssim\lambda_{r}$ and condition (23) holds. Let $d_{\max}=\max_{k}d_{k}$. Then, after at most $M=O(\log(\psi_{0}/\psi^{\text{\footnotesize ideal}}))$ iterations of Algorithm 3, with probability at least $1-T^{-C}-d^{-C}$, the final estimator satisfies
$w_{i}^{-1}\left|\widehat{w}_{i}\widehat{f}_{it}-w_{i}f_{it}\right|\leq C\left(\sqrt{\frac{d_{\max}\log d}{\lambda_{r}T}}+\sqrt{\frac{1}{\lambda_{r}}}\right)$ (26)
for $1\leq i\leq r,1\leq t\leq T$, and some constant $C>0$.
Note that $w_{i}$ is a scalar representing the factor strength. In a strong factor model, consistent with Bai, (2003); Stock and Watson, (2002); Chen and Fan, (2023), it is typically set as $w_{i}=\sqrt{d}=\sqrt{d_{1}d_{2}\cdots d_{K}}$, and we use $\widehat{w}_{i}\widehat{f}_{it}/w_{i}$ as the estimated factors for further interpretation and analysis.
@@
[…]
where $P_{a_{ik},\perp}=I_{d_{k}}-a_{ik}a_{ik}^{\top}$ and $\Theta=(\Theta_{ij})_{r\times r}$, with $\Theta$ defined in Assumption 4.2. This enables the determination of asymptotic distributions for linear forms of $a_{ik}$.
The following theorem shows the asymptotic distribution of a linear form of the factor loading vector $u^{\top}a_{ik}$ for some fixed vector $u$. Note that in the strong factor model, we have $\Theta_{ii}\asymp w_{i}^{2}\asymp d$ for all $1\leq i\leq r$.
@@
Suppose the conditions in Theorem 4.2 are satisfied. Let $\lambda_{1}\asymp\lambda_{r}\asymp\Theta_{ii}$. Assume that $\lim\inf_{d_{k}\to\infty}$ $\|P_{a_{ik},\perp}u\|_{2}>0$, for each $1\leq i\leq r,1\leq k\leq K$, we have:
@@
    If $T/(d_{k}\lambda_{r})\to 0$, then
    $\sqrt{T\Theta_{ii}}\sigma_{u,ik}^{-1}u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)\xrightarrow{d}N(0,1),$ (27)
    where $\sigma_{u,ik}^{2}=h_{ik}^{\top}\Sigma_{e}h_{ik}$, $h_{ik}=b_{iK}\odot\cdots\odot b_{i,k+1}\odot P_{a_{ik},\perp}u\odot b_{i,k-1}\odot\cdots\odot b_{i1}\in\mathbb{R}^{d}$ and $\odot$ represents Kronecker product.
  2. (ii)
    If $d_{k}\lambda_{r}=O(T)$, then
Remark 4.5.
Theorem 4.4 is analogous to Theorem 2 of Bai, (2003). The dominant case is (i), which exhibits asymptotic normality, while case (ii) is of theoretical interest when a specific convergence rate is required. In the typical strong factor model, we have $\Theta_{ii}\asymp w_{i}^{2}\asymp d$ for all $1\leq i\leq r$. Thus if $T/(d_{k}d)\rightarrow 0$, case (i) applies; conversely, if $d_{k}d=O(T)$, the error from the noise covariance dominates, leading to case (ii). In practice, we can estimate $\Theta$ by $\widehat{\Theta}=\frac{1}{T}\sum_{t=1}^{T}\widehat{W}\widehat{F}_{t}\widehat{F}_{t}^{\top}\widehat{W}$ and then compare $d_{k}\widehat{\Theta}_{ii}$ with $T$ to determine the appropriate case. In the special case where the noise tensor ${\cal E}_{t}$ has i.i.d. entries, only case (i) is applicable.
Remark 4.6.
The asymptotic normality results provide theoretical guarantees that allow us to construct confidence intervals and conduct hypothesis tests. For example, with consistent estimators $\widehat{\Theta}$ and $\widehat{\sigma}_{u,ik}^{2}=\widehat{h}_{ik}^{\top}\widehat{\Sigma}_{e}\widehat{h}_{ik}$, we can form a $(1-\alpha)$-level confidence interval for $\widehat{a}_{ijk}^{{\rm iso}}$ as $\left(\widehat{a}_{ijk}^{{\rm iso}}-q_{1-\alpha/2}\sqrt{T\widehat{\Theta}_{ii}}\widehat{\sigma}_{u,ik}^{-1},\right$. $\left.\widehat{a}_{ijk}^{{\rm iso}}+q_{1-\alpha/2}\sqrt{T\widehat{\Theta}_{ii}}\widehat{\sigma}_{u,ik}^{-1}\right)$, $1\leq i\leq r,1\leq j\leq d_{k},1\leq k\leq K$, where $q_{1-\alpha/2}$ is the $1-\alpha/2$ quantile of the standard normal distribution. Figures 11 and 13 in Section 6 illustrate this by reporting estimated loadings for each characteristic and decile with 95% confidence intervals. Beyond this, asymptotic normality opens the door to a range of testing strategies that are central in empirical asset pricing. As Bai, (2003) emphasizes, estimated factor loadings can be used to construct portfolios in the spirit of Lehmann and Modest, (1998). Similarly, Lettau, (2024) interpret Tucker factor loadings as portfolio weights. Our asymptotic distribution theory for CP loadings makes analogous portfolio-based inference feasible: for example, comparing the relative importance of portfolio weights associated with different characteristics [^11] A further illustration comes from alpha testing. In empirical asset pricing, an important question is whether portfolio alphas vanish after controlling for observed and latent factors, as in Giglio et al., (2021). Our CP factor model could potentially be extended to this context. The asymptotic normality results in Theorem 4.4 provide the foundation for constructing valid t-statistics for such tests. Exploring this extension would require substantial additional work and therefore lies beyond the scope of the present paper, but it illustrates how our inferential framework can, in principle, be applied to important empirical questions in asset pricing..
Since $\sigma_{u,ik}^{2}\asymp d_{k}$, Theorem 4.4(i) implies that $u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)=O_{\mathbb{P}}(\sqrt{d_{k}/(\lambda_{r}T})$. In contrast, Theorem 4.4(ii) yields $u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)=O_{\mathbb{P}}(1/\lambda_{r})$. In Theorem 4.4, we focus on vectors $u$ with the property that $\|P_{a_{ik},\perp}u\|_{2}>0$ when $d_{k}$ is large, which effectively assumes $\sin\angle(u,a_{ik})>0$. Conversely, when $\sin\angle(u,a_{ik})=0$, the convergence rate of the estimated linear form is faster, and its asymptotic distribution is a mixture of $\chi_{1}^{2}$ distributions. [^12] Theorem 4.5 complements Theorem 4.4 by addressing the case where $\sin\angle(u,a_{ik})=0$, i.e., when the direction of interest $u$ is aligned with the loading vector $a_{ik}$. In most economic applications, however, we are interested in inference for a particular component of the loading, in which case $u$ is a unit vector aligned with a coordinate axis. Hence, the condition $\sin\angle(u,a_{ik})>0$ is typically satisfied, and Theorem 4.4 remains applicable. As a safeguard, one can check whether $\|P_{\widehat{a}_{ik},\perp}u\|_{2}$ is close to zero using the estimated $\widehat{a}_{ik}$ and the user-specified $u$. Similarly, we will establish that
@@
$+O_{\mathbb{P}}\left(\frac{1}{\Theta_{ii}^{2}}\left(\frac{d_{k}\log(d)}{T}+\sqrt{\frac{d_{k}\log(d)}{T}}+1\right)^{2}\right).$
@@
Suppose the conditions in Theorem 4.2 are satisfied. Let $\lambda_{1}\asymp\lambda_{r}\asymp\Theta_{ii}$. For each $1\leq i\leq r,1\leq k\leq K$, we have:
@@
    If $T/(d_{k}\lambda_{r})\to 0$, then
    $T\Theta_{ii}d_{k}^{-1}\left(1-(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})^{2}\right)\xrightarrow{d}\sum_{j=1}^{d_{k}}\varpi_{j}\chi_{1}^{2},$ (30)
    where $\varpi_{j},1\leq j\leq d_{k}$ are the eigenvalues of $\Sigma_{ik}^{1/2}P_{a_{ik},\perp}\Sigma_{ik}^{1/2}$, with $\Sigma_{ik}=d_{k}^{-1}\mathbb{E}[({\cal E}_{t}\times_{\ell\neq k}^{K}b_{i\ell})({\cal E}_{t}\times_{\ell\neq k}^{K}b_{i\ell})^{\top}]$.
@@ 5 Simulation
In this section, we conduct empirical comparisons among different methods for estimating loading vectors across various simulation scenarios, and verify the limiting distribution of the estimated loading vectors. We assess the performance of the contemporary covariance-based iterative simultaneous orthogonalization procedure (CC-ISO) proposed in this paper, auto-covariance-based iterative simultaneous orthogonalization procedure by Han et al., (2024) (AC-ISO), and TPCA by Babii et al., (2022). The auto-covariance considered by Han et al., (2024) is defined by the following lagged-cross product operator:
@@
$\max_{i,k}\|\widehat{a}_{ik}\widehat{a}_{ik}^{\top}-a_{ik}a_{ik}^{\top}\|_{2}=\max_{i,k}\sqrt{1-(\widehat{a}_{ik}^{\top}a_{ik})^{2}}.$
@@
The factor processes $f_{it}$ exhibit weak temporal dependence and are generated as an independent AR(1) process, with the factor strength $w_{i}$ being a scalar depending on $d_{1},d_{2}$ and $r$:
$f_{i,t+1}$ $=\phi f_{it}+\epsilon_{it},$ (32)
$w_{i}$ $=\frac{1}{5}(r-i+1)\sqrt{d_{1}\times d_{2}},$ (33)
where $\phi=0.1$, $\operatorname{Var}(\epsilon)=1-\phi^{2}=0.99$. In this case, the tensor factor model in (1) represents a typical strong factor model, where $\lambda_{i}=w_{i}^{2}=(r-i+1)^{2}d_{1}d_{2}/25=O(d)$ when $r$ is fixed. The results with $\phi=0.5$ are reported in Appendix C.
@@
    (Orthogonal loading matrix) Set $\eta=0$ so that the columns of loading matrix $A_{k}$ are orthonormal. Each entry of error term ${\cal E}_{t}$ is generated independently from $N(0,1)$.
@@
    (Non-orthogonal loading matrix) Vary $\eta$ in the set $\{0.05,0.15,0.25\}$ so that the columns of loading matrix $A_{k}$ are not orthogonal. Each entry of error term ${\cal E}_{t}$ is generated independently from $N(0,1)$.
@@
    (Serial correlation in ${\cal E}_{t}$) Set $\eta=0.1$. The errors ${\cal E}_{t}$ are generated according to ${\cal E}_{t}=\Psi_{1}^{1/2}Z_{t}\Psi_{2}^{1/2}$, where
@@
    (Weak factors) Set $r=3$, and $\eta=0.1$. The error terms are generated according to ${\cal E}_{t}=\Psi_{1}^{1/2}Z_{t}\Psi_{2}^{1/2}$, where
@@
    The scaling multiplier in factor process $w_{i}=(r-i+1)(d_{1}d_{2})^{1/\alpha}$, where $\alpha$ varies in the set $\{2.5,3,3.5,4\}$. Note that when $\alpha=2$, the factor structure is considered strong. A larger $\alpha$ indicates a weaker factor structure.
@@
In Configuration II, we assess the impact of non-orthogonal factor loadings on estimation. Figure 3 shows the ratio of the estimation errors of CC-ISO to AC-ISO (first panel) and to TPCA (second panel) across different values of $\eta$ and dimensions. In the first panel, the error ratio of CC-ISO to AC-ISO remains around 0.15, indicating the superior accuracy of CC-ISO. The ratio remains relatively stable because the signal part in AC-ISO, albeit small, also increases with dimensions, limiting its relative improvement. In contrast, in the second panel, the error ratio of CC-ISO to TPCA decreases and converges as dimensions grow. This is because TPCA is unable to recover non-orthogonal factor loading vectors, resulting in stable estimation errors regardless of dimension. Meanwhile, CC-ISO effectively identifies non-orthogonal loadings, leading to estimation errors that diminish and eventually converge to zero.
@@
Figure 4 shows the ratio of estimation errors of CC-ISO to AC-ISO under configuration III, designed to evaluate the robustness of the proposed CC-ISO algorithm against serial correlation in the error term. We observe that CC-ISO’s performance improves monotonically as $T$ increases. In contrast, AC-ISO’s performance deteriorates as the serial correlations in the error term strengthen. This decline is due to the contamination of signals in the auto-covariance by the serial correlations in the error terms. However, CC-ISO demonstrates robustness against such serial correlations.
@@
We also examine the performance of Randomized Projection (RP) from Procedure 2 and compare it with composite PCA.
@@
    (composite PCA vs. RC-PCA) $r=5$. $d_{1}=d_{2}=\bar{d}$ with $\bar{d}\in\{20,40,80\}$ and $T\in\{100,200,500\}$. The columns of factor loadings $A_{k}$ are orthonormal and are generated as described in Configuration I. Furthermore, the factors $f_{it}$ are also orthonormal, generated using QR decomposition after deriving from AR(1) processes. In this setting, the singular values of the common components $\sum_{i=1}^{r}w_{i}f_{it}a_{i1}\otimes a_{i2}$ are solely determined by $w_{i}$. We set $w_{i}=w=10$ to ensure identical eigenvalues of common components. Error terms are generated from i.i.d. $N(0,1)$. Though the top $r$ eigenvalues of the $\widetilde{\Sigma}$ are not identical due to noise, their differences are relatively small, allowing randomized projection algorithms to ensure the accuracy of initial estimations. For the remaining parameters, we set $\nu=0.8$, $c_{0}=0.1$ and $L=2r^{2}$.
@@
Given the close empirical performances of CC-ISO under both initialization methods under configuration V, our focus shifts to the estimation errors of the initial estimations, as illustrated in Figure 6. RC-PCA outperforms composite PCA in terms of the accuracy of initial estimations, particularly pronounced when $d$ is smaller and $T$ is larger. This occurs because the sample covariance of $\operatorname{Vec}({\cal E}_{t})$ approaches the identity matrix when $d$ is small and as $T$ increases. Consequently, $\widetilde{\Sigma}$ is more likely to have eigenvalues that are closer together.
@@
where ${\cal F}_{t}\in\mathbb{R}^{r\times r}$ is the factor process in the Tucker factor model. In the CP factor model, ${\cal F}_{t}$ is diagonal with the $i^{th}$ diagonal element equal to $f_{it}$. In the mis-specification setting, we allow the off-diagonal entries to deviate from 0. Denote the $(i,j)^{th}$ entry of ${\cal F}_{t}$ by ${\cal F}_{t,ij}$. Let ${\cal F}_{t,ij}=w_{ijt}f_{ijt}$, where $f_{ijt}$ is generated as specified in (32). The configuration is as follows:
@@
    (Mis-specification) $r=3,(d_{1},d_{2})\in\{(40,40),(40,60),(60,60)\}$ and $T\in\{100,300,500\}$. The loading vectors and error terms are generated as in Configuration IV, allowing for correlation between loading vectors and weak cross-sectional correlation in the error term. Set $w_{iit}=\sqrt{d_{1}d_{2}}/5$ and $w_{ijt}=(d_{1}d_{2})^{1/\alpha}/5$ if $i\neq j$ with $\alpha\in\{3,4,5\}$. A smaller $\alpha$ indicates a more severe mis-specification in the model.